OpenStudy (anonymous):

Use substitution to solve the system of equations. x=2y-5 x+3y=10

OpenStudy (marissalovescats):

Oh by substitution okay ' So it says: x=2y-5 This means we plug that in for x in the second equation

OpenStudy (marissalovescats):

2y-5+3y=10 add 5 to both sides: 2y+3y=15 2y+3y is 5y so 5y=15 divide by 5 y=3

OpenStudy (marissalovescats):

Since we know y=3 we plug that into: x=2y-5 x=2(3)-5 x=6-5 x=1 So we solved it. x=1 and y=1

OpenStudy (marissalovescats):

y=3* oops typo :P

OpenStudy (anonymous):

So because there isn't a multiplication option, you would just add the x equation to the other equation and not multiply? That confuses me sometimes.

OpenStudy (marissalovescats):

No you don't have to do any multiplication. Substitution just means put it in place of. So since we were given x=2y-5 this means that is what x equals. so in the 2nd equation we just replace x with 2y-5

OpenStudy (anonymous):

Thank you :) This helped a lot.

OpenStudy (marissalovescats):

You're welcome :)